tan(37.5)=a^1/2-b^1/2+c^1/2-d^1/2 wher a,b,c,d belongs to N and a>b>c>d then ad/bc is equal to 
 

Grade:11

1 Answers

Subash kumar
5 years ago
Tan(37.5)=Tan ( 45-7.5)=(1-Tan 7.5)/(1+tan 7.5)=(cos 7.5 - sin 7.5)/(cos 7.5+ sin 7.5)=(1-sin 15)/cos 15, now put sin 15=(√3-1)/2√2 and  cos 15=(√3+1)/2√2 tan 37.5=√6+√3-√4-√2 so ac=6*3=18 and bd=4*2=8,ac/bd=18/8=9/4